OpenStudy (anonymous):

what is the mass of an object with a density of 5.7g/cm3 and a volume of 48ml?

OpenStudy (anonymous):

d=m/V m=Vd=48*5.7 grams

OpenStudy (anonymous):

273.6

OpenStudy (anonymous):

but what about the units

OpenStudy (akashdeepdeb):

grams !! :)